Service Status

ADSL/FTTC/FTTP Connectivity:

Ethernet/Leased Line Connectivity:

HVS Cloud Telephone Services:

Cloud VPS/Hosting Servers:

Scheduled Maintenance:

 

Further information:

ALL RUNNING AS EXPECTED.

Outage at CID Leeds due to JCB having cut through fibre optic cables.

ALL RUNNING AS EXPECTED.

ALL RUNNING AS EXPECTED.

NO MAINTENANCE SCHEDULED.

12/10 @ 5pm – Engineers are due to attend site tomorrow but due to the complexity of re-routing and repairing damaged cable we currently have no estimate of completion timeframe. Further update due 13/10 @ 11am.

13/10 @ 10:30am – Engineers attended site at 8am this morning. Discovered completely damaged duct work and shredded fibre in the dig hole which means a full action plan is required prior to being able to facilitate repairs. Next update due @ circa 1:30pm.

13/10 @ 2:00pm – Engineers are still working on an action plan with input from Leeds CC planning due to location of damage. Further update due around 4pm.

Menu